Two numbers added together equal 40. Twice the first number is 3 more than 5 times the other number. What are the two numbers?
Answer by stanbon(57203)

Two numbers added together equal 40. Twice the first number is 3 more than 5 times the other number. What are the two numbers?
----
Equations:
x + y = 40
2x= 5y+3
----------------
Substitute for "x" and solve for "y":
2(40-y) = 5y+3
80-2y = 5y+3
7y = 77
y = 11
----
Solve for "x":
x + y = 40
x + 11 = 40
x = 29
==================
Cheers,
Stan H.